Peter's brother, clement is seven years younger than he is. Twenty years ago, Clement's age was only three-quarters of Peter's age. How old is Clement now
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

Peter's brother, clement is seven years younger than he is. Twenty years ago, Clement's age was only three-quarters of Peter's age. How old is Clement now
----------------------
P = C - 7
P-20 = (3/4)*(C - 20)
==========
C-7 - 20 = (3/4)*(C-20)
C - 27 = (3/4)*(C-20)
4C - 108 = 3C - 60
C = 48

Answer by greenestamps(13200)



The difference in their ages -- now and at any other time -- is 7 years.

So when Clement was 3/4 as old as Peter, their ages were 21 and 28.

(If you don't see that with mental calculations, here is the algebra:
Clement's age = 3x
Peter's age = 4x
difference = x = 7
Clement = 21; Peter = 28)

That was 20 years ago, so now their ages are 41 and 48.

ANSWER: Clement is now 41